OpenStudy (anonymous):

(6-i)/1+i ..how to simplify it . what rule applies here>??

OpenStudy (helder_edwin):

\[ \large \frac{6-i}{1+i}\cdot\frac{1-i}{1-i}= \]

OpenStudy (anonymous):

I did this step . then I am confused in the next step

OpenStudy (anonymous):

Which part u confused?

OpenStudy (cwrw238):

multiply top and bottom by the conjugate of 1+i = 1 - i

OpenStudy (anonymous):

just times it out like how u multiply quadratics

OpenStudy (anonymous):

so the bottom is not 1-i^2?

OpenStudy (anonymous):

(a+b)(a+b) << same as (1+i)(1-i)

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

i^2 is equivalent to negative 1

OpenStudy (anonymous):

so it will be 1 - (-1)

OpenStudy (anonymous):

and what about the denumerator ? in the answer it is 2?

OpenStudy (anonymous):

numerator is 5-7i

OpenStudy (anonymous):

so 1-i and 1-i add?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(6-i)(1-i)= (6 - 6i -1 + i^2)

OpenStudy (anonymous):

yes this is what I got . I am confused about the denumerator

OpenStudy (anonymous):

how to multiply it?

OpenStudy (anonymous):

denumerator...? u mean denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(1+i)(1-i) (1 -i + i -i^2)

OpenStudy (anonymous):

because i^2 = -1

OpenStudy (anonymous):

so it will be (1 - (-1) ) which equals 2

OpenStudy (anonymous):

ohh so I have to solve it too by rule

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

very help ful . its was not mentioned in the book .. Thanx

OpenStudy (anonymous):

no problem xD by the way, i= square root of -1 i^2 is just 1 i^3 = i and so on...